# Graph Topology Estimation of Power Grids using Pairwise Mutual Information of Time Series Data

### v2

@@ Abstract
The topology of a power grid is estimated using an information theoretic approach. By modeling the grid as a graph and using voltage magnitude data of individual nodes in the grid, the mutual information between pairs of nodes is computed using different approximation methods. Using the well-known Chow-Liu algorithm, a maximum spanning tree based on mutual information is computed to estimate the power grid topology. This manuscript explores the application of this method to different datasets and explores the domain of applicability. The data quality, precision, time windows, frequency and the method for calculating the mutual information are varied to see the effect on the successful reconstruction of the graph and it’s leaf nodes. Success shown for IEEE networks generated with MATPOWER and data generated using GridLAB-D. The algorithm is then cross-validated on IEEE networks.
@@ 1 Introduction
New distributed energy resources are gaining widespread use in the electrical grids of many developed and less-economically developed countries. Solar and wind power plants cause daily reconfiguration in the electrical grid topology, controlled often by a vast array of digitally controlled switches. For the ever-changing topology of the grid, options to save energy, reduce costs and increase load stability are all very important tasks. The true topology of the electrical grid is often a useful tool in many simulations and control algorithms to optimize power transmission Liao et al. (2018).
@@
The grid utility companies often possess outdated or mislabeled electrical grid maps Park et al. (2020). An untrue map impairs the ability of the utility company to observe power flow and optimize power flow. For instance, the utility company might want to decide where to place a 500 kWh flow-battery to improve load stability. Knowing the topology of the grid can help the utility company optimize placement of the battery. Often the utility companies will have a partial topology of the electrical grid which has been confirmed Lampe and Ahmed (2013), and they wish to use this partial topology to confirm whether outlying sub-sections of the grid are correctly mapped.
@@
Previously, researchers in Ref. Liao et al. (2015, 2016) have defined the problem statement for the estimation of the graph structure of a power grid using smart-meter data. We can define a distribution grid as a graph with edges (branches) that connect different nodes (buses). To utilize the time series data collected by smart meters, for a $M$ bus system, we construct a graphical model $G=(S,E)$ where $S=\{1,2,...M\}$ is the set of vertices, and $E=\{e_{ik},i,k\in S\}$ represents the set of the unidirectional edges meaning $e_{ik}=e_{ki}$. In this model, which is common in the literature Poudel et al. (2023); Zhu et al. (2024); Zhao et al. (2025) a node corresponds to a bus in the physical layer and is modeled as a random variable $V_{i}$. The edge that connects node $i$ and $k$ represents the statistical dependence measured between bus $i$ and $k$. If we assume for a noiseless system that measurements are taken in steady state and all voltages are sinusoidal signals at the same frequency, then the voltage measurement at bus $i$ is defined as $v_{i}[t]=\lvert v_{i}[t]\rvert e^{(j\delta_{i}[t])}$, where $\lvert v_{i}[t]\rvert$ is the voltage magnitude of node $i$ at time $t$ and $\delta_{i}[t]$ is the voltage phase at node $i$ at time $t$. This works seeks to reconstruct the grid bus topology using voltage magnitude data and phase data when available. In other words, given a sequence of of historical measurements $\lvert v_{i}(t)\rvert$ we seek to find graph structure of the power grid. In some datasets we will also have access to sensor current data (magnitude and phase, $\delta_{i}[t]$ and i[t]{i}[t] respectively) as well.
@@
Previous research Weng et al. (2016) showed success using the pairwise mutual information of sensor-data and using the Chow-Liu algorithm Chow and Liu (1968); Koller and Friedman (2009) to reconstruct the power grid structure Weng et al. (2016). In this work, we extend this method to different methods of mutual information approximation and test the limits of this method on new data while varying the length of data (in days), the resolution of the smart-meter data (in bits and rounding to a significant digit), the downsampling rate. This experiments are important for operators to understand what length and precision of data are needed to store Hernández et al. (2025); Engel et al. (2025). We also test different entropy approximations which speed up the method’s computation time of the mutual information Chen and Wang (2021). We present the following contributions.
@@ 2.1 Theoretical estimation approach
Voltage measurements usually have irregular probability distributions. Instead of using voltage measurements directly, we use the incremental change of the voltage measurement to reconstruct the distribution grid topology Chen et al. (2016). A reasonable assumption is that the change in current at bus $i$ is independent given the change in current at neighboring nodes. Given this assumption, the change in voltage distribution at a bus is conditionally independent to the change in voltage distributions of non-neighboring buses given the change in voltage distributions of neighboring buses Liao et al. (2016). Figure 16 in the Appendix shows pairwise mutual information of the change in current phasors. The relatively small mutual information means that one can approximate that the current are independent with some approximation error.
@@
Previous work found that the pairwise mutual information based methods to found that an maximum spanning tree algorithm with mutual information as weights optimally approximates $P(V_{S})$ in Equation (1Weng et al. (2016). We re-state the theorem from this work below without proof in Theorem 1.
@@
With these theorems in mind, we employ the Chow-Liu algorithm Chow and Liu (1968) for our topology estimation problem. The algorithm finds the maximum weight spanning tree based on mutual information between nodes. The mutual information is defined as in Equation (4) for continuous random variables.
@@
5:Sort all possible bus pair $(i,j)$ into non-increasing order by $I(V_{i},V_{j})$. Let $\tilde{\mathcal{E}}$ denote the sorted set.
6:Let $\hat{\mathcal{E}}$ be the set of nodal pair comprising the maximum weight spanning tree. Set $\hat{\mathcal{E}}$$\emptyset$
@@
16:  return $\hat{\mathcal{E}}$
17:end for
We can visualize the algorithm steps in Figure 1 as a data analysis workflow (DAW) Liew et al. (2016); Schintke et al. (2024). If we can accurately compute the mutual information between nodes, we can run the Chow-Liu algorithm to return an estimate of the electrical grid topology. Our solution therefore relies on calculating the mutual information between a pair of voltage distributions at distinct nodes.
@@
We can further visualize how mutual information of pairs of nodes determines the branches we estimate. In Figure 2, we can see the heatmap of mutual information for a 34 bus network. The self-information terms Stone (2024), $MI(V_{i},V_{i})$, have the largest values in the heatmap but these values are not included in our calculations since in the Chow-Liu algorithm we look at the mutual information for pairs of distinct nodes. The heatmap shows that many of the largest MI values in the heatmap are marked with a green cross to show that the algorithm has estimated an edge between these two nodes.
@@ 2.2 Gaussian approximation
Our program must therefore accurately compute the mutual information between voltage distributions at distinct pairs of nodes. The mutual information can be reformatted in terms of entropy terms as in Equation (5) Cover and Thomas (2012).
@@
The first method to compute mutual information is a Gaussian approximation method. To understand the method, we note that that the entropy of a discrete random variable $x$ is defined as in Equation (6Lemons (2013).
@@
$H(X)=\frac{k}{2}(1+\ln{2\pi})+\frac{1}{2}\det\Sigma$ (7)
@@ 2.3 Discrete mutual information
As mentioned earlier, the more computationally expensive option for large data sets is to approximate probabilities for frequencies of occurrence in the data sets. The mutual information can also be calculated as in Equation (8). The discrete estimate should yield the most accurate mutual information for large data sets since the frequency of observing a discrete value approaches the true probability for the discrete value MacKay (2003). Note this process is quite slow for a network where nodes can take on many values and there is a large amount of data. Thus, we expect the discrete mutual information to give us the most accurate value but with the slowest performance.
@@ 2.4 JVHW Entropy Estimation
Due to the computational expense of calculating the discrete MI of data, many research groups have explored computational short-cuts Walters-Williams and Li (2009). The Weissman group at Stanford University has developed minimax estimators to optimally estimate the entropy of a function Jiao et al. (2015). The estimators treat the smooth and non-smooth region of the function differently. A bias corrected MLE is applied to the smooth region of the function and the best polynomial fit to the non-smooth region. The JVHW estimator seeks to estimate a functional of entropy, $F_{\alpha}(P)$ defined as:
@@
where $\alpha>0$ for a unknown discrete probability distribution $P=(p_{1},p_{2},...,p_{S})$, with unknown support size $S$. Regarding estimation of the the functional of entropy, the $L_{2}$ risk of an arbitrary estimator, $\hat{F}$, is defined as:
@@ 2.5 Data sets
The data used in this article was generated using MATPOWER Zimmerman et al. (2010) a well-known grid simulation package for Matlab. The work described in this document seeks to reproduce the results of the Rajagopal group for an IEEE 123-bus network data generated with MATPOWER and extend results to data generated with GridLAB-D Chassin et al. (2008), a software package developed by the Pacific Northwest National Laboratory to simulate power flow and control of electrical grids. This manuscript presents the following listed contributions.
@@
In the IEEE data sets, we have one year hourly yearly voltage magnitude and phase data for each bus. Hence, there are $8760$ data points in the IEEE data set for each bus. Similarly, the GridLAB-D data sets cover a year of data but have more precise resolution with a measurement taken every minute. This gives $525,610$ data points for each bus.
@@ 3.2 Gaussian Approximation of Data
To get a better sense of the nature of the different data sets, we may examine the Kullback-Leibler (KL) divergence Kullback and Leibler (1951) between the voltage magnitude at one node and the a Gaussian fit based on the data’s mean and variance. The KL divergence between two probability distributions, $p$ and $q$, is a measure of similarity of the distributions. The formula for the KL divergence is shown in equation (14. Note, unlike mutual information, the KL divergence is not symmetric.
@@
To compute the KL-divergence the data at each node are binned into a thousand different bins. The Gaussian approximation for the data is also computed at the same thousand voltage magnitude (or change in voltage magnitude) values as the binned data. The KL-divergence is then computed by using both sets of data and equation (14). The histogram and Gaussian approximation for the change in voltage magnitude at Node 33 of the SG1 data set is shown in the Appendix in Figure 19.
The mean, min, max, and standard deviation of the KL divergence between the change in voltage magnitude histogram and the Gaussian approximation for each data set is shown in Table (LABEL:tb:mag_KL). Note, the KL divergence is computed at each node in the data set and the mean refers to the mean KL divergence across all nodes in a specific data set. The SG1 non-solar and SG2 solar/non-solar data sets all have similar mean KL divergence at around 7 bits. The SG1-solar shows a slightly worse fit at above 8 bits.
@@ 3.3 Reducing Data Requirements for Estimation
Most smart meters deployed in Canada sample at the most frequent rate of 15 minutes Lee and Hess (2021). In California older models sampling once every half hour or once every hour in residential buildings Li et al. (2021). In Norway, the UK and France, this sampling happens every thirty minutes Lee and Hess (2021). We down-sample the data sets generated in GridLAB-D from one sample every 1 minute to commercially available resolutions such as 5, 15, 30, or 60 minutes per data point and observe the effect on the SDR.
@@
Moreover, the SG2 (solar) data shows that that the JVHW method improves slightly at 5 min resolution compared to 1 min, which goes against our expectations for lower SDR with lower resolution. It’s also odd that the SDR for the JVHW method improves at 30 min resolution compared to 15 min resolution. The Gaussian method is the best at 1 min resolution. The Gaussian method has less than a 0.5% percent greater SDR at 1 min res compared to discrete and approximately 2% percent improvement at 5 min resolution but no improvement compared to the discrete method at 15 min resolution. The Gaussian method at 60 min resolution shows the best leaf SDR in Figure 25. We see an approximately 2% increase from 1 minute to 60 minute resolution for the Gaussian method. The leaf node SDR is more than 7% lower than the regular SDR for all MI methods for the SG2 (solar) data set. This suggests the great majority of estimation mistakes for the grid topology are made in estimating leaf branches.
@@ 3.3.1 Data Length Analysis
The effect on the leaf SDR from length size is shown in the Appendix Figure 27 for the SG1 (non-solar) data set. The JVHW and discrete MI methods outperform the Gaussian method. The leaf node mean SDR is greater than 99.5% for the discrete MI methods for a length of data of seven days. A similar plot is show for the SG1 (solar) data set in the Appendix Figure 28. Less data, a 4 day length, are required for greater than 99.5% SDR for the discrete MI methods for the SG1 (solar) data set compared to the non-solar data.
The effect on the SDR of different length sizes of data is shown for the SG2 (non-solar) data set in the Appendix Figure 29. With only two weeks of data all MI methods are less than 0.8% away from the full years SDR. The Gaussian method performs the best of all MI methods for all length sizes. The Gaussian method is within 1% of the full length of data with only one day’s worth of data (at 1 min resolution).
@@
The SG2 solar/non-solar data sets show a rather stagnant SDR from a 30 day length of data to a full year. We expect the SDR to increase with a larger data set. Of interest is whether the same nodes are continuously wrongly predicted as the length of data is increased. The error frequency is defined to be the frequency that a node is estimated to be connected with an incorrect branch. For the 30 day length of data the error frequency is defined as the average over all of the consecutive lengths. For the 364 day length the error frequency is not averaged since there is only one non-overlapping length in the year long data set. Note, if a node is estimated to have three branches and one is connected incorrectly for all thirty consecutive day lengths, then the error frequency will be equal to one. The error frequency heatmap for the SG2 (solar) data set using the Gaussian mutual information on the change in voltage magnitude is shown in Figure 10. The heatmap shows that the same nodes are consistently connected with a branch. This suggest the algorithm is consistently troubled by the same nodes regardless of data length for the SG2 (solar) incremental change in voltage magnitude data.

### v3

@@ Abstract
The topology of a power grid is estimated using an information theoretic approach. By modeling the grid as a graph and using voltage magnitude data of individual nodes in the grid, the mutual information between pairs of nodes is computed using different approximation methods. Using the well-known Chow-Liu algorithm, a maximum spanning tree based on mutual information is computed to estimate the power grid topology. This manuscript explores the application of this method to different datasets and explores the domain of applicability. The data quality, precision, time windows, frequency and the method for calculating the mutual information are varied to see the effect on the successful reconstruction of the graph and its leaf nodes. Success is shown for IEEE networks generated with MATPOWER and data generated using GridLAB-D. The algorithm is then cross-validated on IEEE networks.
@@ 1 Introduction
New distributed energy resources are gaining widespread use in the electrical grids of many developed and less-economically developed countries. Solar and wind power plants cause daily reconfiguration in the electrical grid topology, controlled often by a vast array of digitally controlled switches. For the ever-changing topology of the grid, options to save energy, reduce costs and increase load stability are all very important tasks. The true topology of the electrical grid is often a useful tool in many simulations and control algorithms to optimize power transmission [liao2018urban].
@@
The grid utility companies often possess outdated or mislabeled electrical grid maps [park2020learning]. An untrue map impairs the ability of the utility company to observe power flow and optimize power flow. For instance, the utility company might want to decide where to place a 500 kWh flow-battery to improve load stability. Knowing the topology of the grid can help the utility company optimize placement of the battery. Often the utility companies will have a partial topology of the electrical grid which has been confirmed [lampe2013power], and they wish to use this partial topology to confirm whether outlying sub-sections of the grid are correctly mapped.
@@
Previously, researchers in Ref. [liao2015distribution, liao2016urban] have defined the problem statement for the estimation of the graph structure of a power grid using smart-meter data. We can define a distribution grid as a graph with edges (branches) that connect different nodes (buses). To utilize the time series data collected by smart meters, for a $M$ bus system, we construct a graphical model $G=(S,E)$ where $S=\{1,2,...M\}$ is the set of vertices, and $E=\{e_{ik},i,k\in S\}$ represents the set of the unidirectional edges meaning $e_{ik}=e_{ki}$. In this model, which is common in the literature [poudel2023topology, zhu2024ggnet, zhao2025power] a node corresponds to a bus in the physical layer and is modeled as a random variable $V_{i}$. The edge that connects node $i$ and $k$ represents the statistical dependence measured between bus $i$ and $k$. If we assume for a noiseless system that measurements are taken in steady state and all voltages are sinusoidal signals at the same frequency, then the voltage measurement at bus $i$ is defined as $v_{i}[t]=\lvert v_{i}[t]\rvert e^{(j\delta_{i}[t])}$, where $\lvert v_{i}[t]\rvert$ is the voltage magnitude of node $i$ at time $t$ and $\delta_{i}[t]$ is the voltage phase at node $i$ at time $t$. This works seeks to reconstruct the grid bus topology using voltage magnitude data and phase data when available. In other words, given a sequence of historical measurements $\lvert v_{i}(t)\rvert$ we seek to find graph structure of the power grid. In some datasets we will also have access to sensor current data (magnitude and phase, $\delta_{i}[t]$ and i[t]{i}[t] respectively) as well.
@@
The literature presents several different approaches to this subject. Graph neural networks have recently become very popular in a wide variety of applied fields from materials science [bechtel2025band, speckhard2025big], chemistry [fung2021benchmarking], social sciences [fan2019graph] and have found success in power grid applications [liao2021review]. These approaches are, however, generally quite data intensive [wu2020comprehensive] and compute intensive despite recent work to minimize the training cost [speckhard2025analysis]. In this work, however, we focus on a specific method [weng2016distributed], namely using the pairwise mutual information of sensor-data and using the Chow-Liu algorithm [chow1968approximating, koller2009probabilistic] to reconstruct the power grid structure [weng2016distributed]. In this work, we extend this method to different methods of mutual information approximation and test the limits of this method on new data while varying the length of data (in days), the resolution of the smart-meter data (in bits and rounding to a significant digit), the downsampling rate. These experiments are important for operators to understand what length and precision of data are needed to store [hernandez2025monitoring, engel2025real]. We also test different entropy approximations which speed up the method’s computation time of the mutual information [chen2021efficient]. We present the following contributions [speckhard2025grid].
@@ 2.1 Theoretical estimation approach
Voltage measurements usually have irregular probability distributions. Instead of using voltage measurements directly, we use the incremental change of the voltage measurement to reconstruct the distribution grid topology [chen2016quickest]. A reasonable assumption is that the change in current at bus $i$ is independent given the change in current at neighboring nodes. Given this assumption, the change in voltage distribution at a bus is conditionally independent to the change in voltage distributions of non-neighboring buses given the change in voltage distributions of neighboring buses [liao2016urban]. Figure 16 in the Appendix shows pairwise mutual information of the change in current phasors. The relatively small mutual information means that one can approximate that the current are independent with some approximation error.
@@
Previous work found that the pairwise mutual information based methods to find that a maximum spanning tree algorithm with mutual information as weights optimally approximates $P(V_{S})$ in Equation (1[weng2016distributed]. We re-state the theorem from this work below without proof in Theorem 1.
@@
With these theorems in mind, we employ the Chow-Liu algorithm [chow1968approximating] for our topology estimation problem. The algorithm finds the maximum weight spanning tree based on mutual information between nodes. The mutual information is defined as in Equation (4) for continuous random variables.
@@
5:Sort all possible bus pairs $(i,j)$ into non-increasing order by $I(V_{i},V_{j})$. Let $\tilde{\mathcal{E}}$ denote the sorted set.
6:Let $\hat{\mathcal{E}}$ be the set of nodal pairs comprising the maximum weight spanning tree. Set $\hat{\mathcal{E}}$$\emptyset$
@@
16:end for
17:return $\hat{\mathcal{E}}$
We can visualize the algorithm steps in Figure 1 as a data analysis workflow (DAW) [liew2016scientific, SCHINTKE202482]. If we can accurately compute the mutual information between nodes, we can run the Chow-Liu algorithm to return an estimate of the electrical grid topology. Our solution therefore relies on calculating the mutual information between a pair of voltage distributions at distinct nodes.
@@
We can further visualize how mutual information of pairs of nodes determines the branches we estimate. In Figure 2, we can see the heatmap of mutual information for a 34 bus network. The self-information terms [stone2024information], $MI(V_{i},V_{i})$, have the largest values in the heatmap but these values are not included in our calculations since in the Chow-Liu algorithm we look at the mutual information for pairs of distinct nodes. The heatmap shows that many of the largest MI values in the heatmap are marked with a green cross to show that the algorithm has estimated an edge between these two nodes.
@@ 2.2 Gaussian approximation
Our program must therefore accurately compute the mutual information between voltage distributions at distinct pairs of nodes. The mutual information can be reformatted in terms of entropy terms as in Equation (5) [cover2012elements].
@@
The first method to compute mutual information is a Gaussian approximation method. To understand the method, we note that that the entropy of a discrete random variable $x$ is defined as in Equation (6[lemons2013student].
@@
$H(X)=\frac{k}{2}(1+\ln{2\pi})+\frac{1}{2}ln(\det\Sigma)$ (7)
@@ 2.3 Discrete mutual information
As mentioned earlier, the more computationally expensive option for large datasets is to approximate probabilities for frequencies of occurrence in the datasets. The mutual information can also be calculated as in Equation (8). The discrete estimate should yield the most accurate mutual information for large datasets since the frequency of observing a discrete value approaches the true probability for the discrete value [mackay2003information]. Note this process is quite slow for a network where nodes can take on many values and there is a large amount of data. Thus, we expect the discrete mutual information to give us the most accurate value but with the slowest performance.
@@ 2.4 JVHW Entropy Estimation
Due to the computational expense of calculating the discrete MI of data, many research groups have explored computational short-cuts [walters2009estimation]. The Weissman group at Stanford University has developed minimax estimators to optimally estimate the entropy of a function [jiao2015minimax]. The estimators treat the smooth and non-smooth region of the function differently. A bias corrected MLE is applied to the smooth region of the function and the best polynomial fit to the non-smooth region. The JVHW estimator seeks to estimate a functional of entropy, $F_{\alpha}(P)$ defined as:
@@
where $\alpha>0$ for a unknown discrete probability distribution $P=(p_{1},p_{2},...,p_{S})$, with unknown support size $S$. Regarding estimation of the functional of entropy, the $L_{2}$ risk of an arbitrary estimator, $\hat{F}$, is defined as:
@@ 2.5 Datasets
The data used in this article was generated using MATPOWER [zimmerman2010matpower], a well-known grid simulation package for Matlab. The work described in this document seeks to reproduce the results of the Rajagopal group for an IEEE 123-bus network data generated with MATPOWER and extend results to data generated with GridLAB-D [chassin2008gridlab], a software package developed by the Pacific Northwest National Laboratory to simulate power flow and control of electrical grids. This manuscript presents the following listed contributions.
@@
In the IEEE datasets, we have one year of hourly voltage magnitude and phase data for each bus. Hence, there are $8760$ data points in the IEEE dataset for each bus. Similarly, the GridLAB-D datasets cover a year of data but have more precise resolution with a measurement taken every minute. This gives $525,610$ data points for each bus.
@@ 3.2 Gaussian Approximation of Data
To get a better sense of the nature of the different datasets, we may examine the Kullback-Leibler (KL) divergence [kullback1951information] between the voltage magnitude at one node and a Gaussian fit based on the data’s mean and variance. The KL divergence between two probability distributions, $p$ and $q$, is a measure of similarity of the distributions. The formula for the KL divergence is shown in Equation 14. Note, unlike mutual information, the KL divergence is not symmetric.
@@
To compute the KL-divergence the data at each node are binned into a thousand different bins. The Gaussian approximation for the data is also computed at the same thousand voltage magnitude (or change in voltage magnitude) values as the binned data. The KL-divergence is then computed by using both sets of data and Equation 14. The histogram and Gaussian approximation for the change in voltage magnitude at Node 33 of the SG1 dataset is shown in the Appendix in Figure 19.
The mean, min, max, and standard deviation of the KL divergence between the change in voltage magnitude histogram and the Gaussian approximation for each dataset is shown in Table (3). Note, the KL divergence is computed at each node in the dataset and the mean refers to the mean KL divergence across all nodes in a specific dataset. The SG1 non-solar and SG2 solar/non-solar datasets all have similar mean KL divergence at around 7 bits. The SG1-solar shows a slightly worse fit at above 8 bits.
@@ 3.3 Reducing Data Requirements for Estimation
Most smart meters deployed in Canada sample at the most frequent rate of 15 minutes [lee2021data]. In California older models sampling once every half hour or once every hour in residential buildings [li2021characterizing]. In Norway, the UK and France, this sampling happens every thirty minutes [lee2021data]. We down-sample the datasets generated in GridLAB-D from one sample every 1 minute to commercially available resolutions such as 5, 15, 30, or 60 minutes per data point and observe the effect on the SDR.
@@
Moreover, the SG2 (solar) data shows that the JVHW method improves slightly at 5 min resolution compared to 1 min, which goes against our expectations for lower SDR with lower resolution. It’s also odd that the SDR for the JVHW method improves at 30 min resolution compared to 15 min resolution. The Gaussian method is the best at 1 min resolution. The Gaussian method has less than a 0.5% percent greater SDR at 1 min res compared to discrete and approximately 2% percent improvement at 5 min resolution but no improvement compared to the discrete method at 15 min resolution. The Gaussian method at 60 min resolution shows the best leaf SDR in Figure 25. We see an approximately 2% increase from 1 minute to 60 minute resolution for the Gaussian method. The leaf node SDR is more than 7% lower than the regular SDR for all MI methods for the SG2 (solar) dataset. This suggests the great majority of estimation mistakes for the grid topology are made in estimating leaf branches.
@@ 3.3.1 Data Length Analysis
The effect on the leaf SDR from length size is shown in the Appendix Figure 27 for the SG1 (non-solar) dataset. The JVHW and discrete MI methods outperform the Gaussian method. The leaf node mean SDR is greater than 99.5% for the discrete MI methods for a length of data of seven days. A similar plot is shown for the SG1 (solar) dataset in the Appendix Figure 28. Less data, a 4 day length, are required for greater than 99.5% SDR for the discrete MI methods for the SG1 (solar) dataset compared to the non-solar data.
The effect on the SDR of different length sizes of data is shown for the SG2 (non-solar) dataset in the Appendix Figure 29. With only two weeks of data all MI methods are less than 0.8% away from the full year’s SDR. The Gaussian method performs the best of all MI methods for all length sizes. The Gaussian method is within 1% of the full length of data with only one day’s worth of data (at 1 min resolution).
@@
The SG2 solar/non-solar datasets show a rather stagnant SDR from a 30 day length of data to a full year. We expect the SDR to increase with a larger dataset. Of interest is whether the same nodes are continuously wrongly predicted as the length of data is increased. The error frequency is defined to be the frequency that a node is estimated to be connected with an incorrect branch. For the 30 day length of data the error frequency is defined as the average over all of the consecutive lengths. For the 364 day length the error frequency is not averaged since there is only one non-overlapping length in the year long dataset. Note, if a node is estimated to have three branches and one is connected incorrectly for all thirty consecutive day lengths, then the error frequency will be equal to one. The error frequency heatmap for the SG2 (solar) dataset using the Gaussian mutual information on the change in voltage magnitude is shown in Figure 10. The heatmap shows that the same nodes are consistently connected with a branch. This suggests that the algorithm is consistently troubled by the same nodes regardless of data length for the SG2 (solar) incremental change in voltage magnitude data.